i believe its 0 but i just wanted to check my work (:
D Question 13 How many d electrons are in Cr 2. .
Answer #1

CG0q ² chromate let the oxidation of ion car is a x +4(-2) = -2 X x -8 = = -2 +6 Therefore, the oxidation state of cor is +6.